OpenStudy (anonymous):

What polynomial has roots of -4,-2,and 5

OpenStudy (bahrom7893):

when you have a polynomial with roots a, b, c you can just do: (x-a)(x-b)(x-c) = 0

OpenStudy (bahrom7893):

To get the answer.

OpenStudy (anonymous):

Is it x^3-x^2-22x+40

OpenStudy (anonymous):

This is my answer: \(x=-4; x=-2; x=5\) \((x+4)(x+2)(x-5)\) \((x^2+2x+4x+8)(x-5)\) \((x^2+6x+8)(x-5)\) \(x^3+6x^2+8x-5x^2-30x-40\) \(x^3+x^2-22x-40\) Remember that \(x=4\) comes from \(x-4=0\).